fxy(x,y) lo, 3. Let the random variables X and Y have the joint probability density function fxy(x, y) = 0<y<1, 0<x<y otherwiLet the random variable X and Y have the joint probability density function.
Answer #1

= (hraltorf ocykl, o exzy. @ E(XY) xy fxsy (214) dx dy. • hp tarx sit = ]: sy = } ] =ECx) = EŞE (x1y) } f4/3) = 3 * 22.- - EY) = 8 o cual = 1/2 +(x1) = $(,9 + 4) = ECX/4) = 8 x f(*11)ax - 2 = ad olekuE (*) = ELE(x))} 그피ye) = 1 EC4) 글싶들 So, TECH) = / 트 해 et cor (x,y) = EXY) - E() E(Y) 하씀 = 124 soyJeovc x)= 1Note-if there is any understanding problem regarding this please feel free to ask via comment box..thank you